OpenStudy (ny,ny):

HELP an orchestra consists of string instruments and wind instruments in ratio of 9 to 4. There are 91 string and wind instruments in the orchestra. How many string instruments are there? How many wind instruments? Use a system of equations to model each situation.

OpenStudy (lgbasallote):

hmm i see that last answerer wasn't able to help huh

OpenStudy (ny,ny):

mhm, i need to write equations.

OpenStudy (lgbasallote):

actually that last one was pretty close...did you notice that if you increase the total, you're just multiplying the ratio by a number? for example, the total is 13 9*1 + 4*1 = 13 if the total is 26 9*2 + 4*2 = 26 if the total is 39 9*3 + 4*3 = 39 do you see this?

OpenStudy (ny,ny):

yeah.

OpenStudy (ny,ny):

multiplying by 7 will give 91.

OpenStudy (lgbasallote):

good. so if the total si given to be 91...but the "multiplier" is not given, what do you do?

OpenStudy (lgbasallote):

imagine you don't know it's 7...you have to write equations remember?

OpenStudy (lgbasallote):

what do you do in algebra if there is a number you don't know?

OpenStudy (ny,ny):

variable.

OpenStudy (lgbasallote):

right. so you multiply 9 and 4 by variables

OpenStudy (ny,ny):

91=9/4x ?

OpenStudy (lgbasallote):

let's say the variable is x you get 9x + 4x = 91

OpenStudy (lgbasallote):

no..not 9/4 x

OpenStudy (ny,ny):

okay. to solve for x we need the other equation

OpenStudy (lgbasallote):

notice what i was doing a while ago 9*1 + 4*1 = 13 9*2 + 4*2 = 26 9*3 + 4*3 = 39 so if you don't know the multiplier... 9*x + 4*x = 39

OpenStudy (ny,ny):

no wait.

OpenStudy (lgbasallote):

that's 91 not 39

OpenStudy (lgbasallote):

what other equation?

OpenStudy (ny,ny):

"system of equations"

OpenStudy (ny,ny):

i know you can combine 9x and 4x into 13x which will give 7 for x.

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

no need for system of equations

OpenStudy (lgbasallote):

you just have 9x + 4x = 91 <--one equation

OpenStudy (lgbasallote):

the rest are just substitutions

OpenStudy (lgbasallote):

i don't know if you call that a system...but it's substitution

OpenStudy (lgbasallote):

to get the number of wind, do 9x to get the number of string, do 4x

OpenStudy (lgbasallote):

or maybe...this is what you mean w = 9x s = 4x w + s = 91 9x + 4x = 91

OpenStudy (lgbasallote):

i have to leave now...so i hope you get your answer...

OpenStudy (ny,ny):

63 string, 28 wind.

OpenStudy (ny,ny):

w+s=91 w=9x s=4x ------ 9x+4x=91 9(7)+4(7)=91 63+28=91 91=91

OpenStudy (ny,ny):

yeah i got it. thank you :D